OpenStudy (anonymous):

A system of linear inequalities is shown below: y ≥ -2x - 2 y + x ≥ 0 Describe the steps to graph the solution set to the system of inequalities.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

to do these problems, you need to be able to graph a line given its equation for example, can you graph the line y= -2x-2 ?

OpenStudy (anonymous):

yes i can graph using geogebra

OpenStudy (phi):

it might help if you graph it.

OpenStudy (anonymous):

i graphed it but i don't see how it would help with getting the steps to do it

OpenStudy (phi):

next, you look at the relation y ≥ -2x - 2 as you know , there are two sides the line. we want the side "above" the line. if that is not clear, I choose a simple point like (0,0) does (0,0) make the relation true ? any idea ?

OpenStudy (phi):

refresh your browser if that did not display properly

OpenStudy (anonymous):

sorry something happened

OpenStudy (anonymous):

and i don't understand it

OpenStudy (phi):

\[ y \ge -2x -2 \] does (0,0) make the relation true ? in other words, with x=0 and y=0: \[ 0 \ge -2 \cdot 0 - 2 \] is this true ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i think

OpenStudy (phi):

simplify the right side (and pay attention to order of operations, multiply before add/subtract)

OpenStudy (anonymous):

yes its true

OpenStudy (phi):

what did you get for the right side ?

OpenStudy (anonymous):

-2

OpenStudy (phi):

ok, and yes 0 is bigger than -2, so the relation is true. all of that means is that the point (0,0) is the "correct side" of the line shade in the area on the side of the line that contains (0,0) so far, so good?

OpenStudy (anonymous):

umm sorry again openstudy is giving me problems and yes so far so good

OpenStudy (anonymous):

@phi so what next?